OpenStudy (anonymous):

2. Let g(x) be the reflection of f(x) = x² + 5 in the x­axis. What is a function rule for g(x)? (1 point) g(x) = –x² – 5 g(x) = x² + 5 g(x) = x² – 5 g(x) = –x² + 5

OpenStudy (anonymous):

@kourtney2014

OpenStudy (anonymous):

g(x)=-x^2-5

OpenStudy (anonymous):

comment all the questions

OpenStudy (anonymous):

oh.. ok

OpenStudy (anonymous):

1. If a function f(x) is shifted to the left one unit, what function represents the transformation? (1 point) f(x + 1) f(x – 1) f(x) + 1 f(x) – 1 2. Let g(x) be the reflection of f(x) = x² + 5 in the x­axis. What is a function rule for g(x)? (1 point) g(x) = –x² – 5 g(x) = x² + 5 g(x) = x² – 5 g(x) = –x² + 5 3. What transformations change the graph of f(x) to the graph of g(x)? f(x) = –7x² g(x) = –35x² + 5 (1 point) The graph of g(x) is the graph of f(x) stretched vertically by a factor of and translated down 5 units. The graph of g(x) is the graph of f(x) stretched vertically by a factor of 5 and translated down 5 units. The graph of g(x) is the graph of f(x) stretched vertically by a factor of and translated up 5 units. The graph of g(x) is the graph of f(x) stretched vertically by a factor of 5 and translated up 5 units. 4. Write an equation for the following transformation of y = x: a vertical stretch by a factor of 6. (1 point) 5. What transformations change the graph of to the graph of ? (1 point) 5. What transformations change the graph of to the graph of ? (1 point) The graph of g is the graph of f translated to the right 3 units and up 7 units. The graph of g is the graph of f translated to the left 3 units and down 7 units. The graph of g is the graph of f translated up 3 units and to the left 7 units. The graph of g is the graph of f translated down 3 units and to the right 7 units. 6. What is the graph of the inequality? 3x + 2y > –5 (1 point) 7. Write an inequality for the graph. (1 point) –6x + 5y ≥ –30 –6x + 5y ≤ –30 5x – 6y ≤ –30 5x – 6y ≥ –30 8. What is the graph of the absolute value equation? y = |x – 5| – 4 (1 point)

OpenStudy (anonymous):

some of it didnt register... like the graphs

OpenStudy (anonymous):

1. f(x+1) 2. g(x)=-x^2+5 3. the graph of g(x) is the graph is the graph of f(x) stretched vertically by a factor of 5 and translated up 5 units. 4. i dont know sorry 5. dont know sorry

OpenStudy (anonymous):

do you have anymore???

OpenStudy (anonymous):

ok. do u have the others?

OpenStudy (anonymous):

there is 8 questions

OpenStudy (anonymous):

is there any more

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

did u see the other above?

OpenStudy (anonymous):

ohh i didnt see those,

OpenStudy (anonymous):

@kourtney2014 im online

OpenStudy (anonymous):

@q12157 can you help me with the questions above? some of them i cant get the pics so just forget those

OpenStudy (anonymous):

@nincompoop @Teddyiswatshecallsme

OpenStudy (anonymous):

I can help if no one helps you in maybe 2-3 hours. I have a 5 page essay due in 20 min in class. Pulled an all nighter.

OpenStudy (anonymous):

ok. thanks. and geez! my sister did that before and stayed awake with monsters lol

OpenStudy (anonymous):

4. y=4x, since y=1/4x would be a vertical compression or horizontal stretch. You can tell because y=4x has a larger slope than y=x 5. Did you copy the question correctly? 6. 3x + 2y > –5 2y>-5-3x therefore y>-2.5-1.5x where y intercept is at (0,-2.5) and x intercept is at (-5/3,0) See below for graph. 7. Write an inequality for the graph. (1 point) –6x + 5y ≥ –30 –6x + 5y ≤ –30 5x – 6y ≤ –30 5x – 6y ≥ –30 What graph? I dont see any graphs... >< 8. What is the graph of the absolute value equation? y = |x – 5| – 4 (1 point) its just horizontally shifted by 5 and vertically shifted by 4. See below for graph.

OpenStudy (anonymous):